# Behavior-Regularized Diffusion Policy Optimization for Offline Reinforcement Learning

### v1

@@ Abstract
The primary focus of offline reinforcement learning (RL) is to manage the risk of hazardous exploitation of out-of-distribution actions. An effective approach to achieve this goal is through behavior regularization, which augments conventional RL objectives by incorporating constraints that enforce the policy to remain close to the behavior policy. Nevertheless, existing literature on behavior-regularized RL primarily focuses on explicit policy parameterizations, such as Gaussian policies. Consequently, it remains unclear how to extend this framework to more advanced policy parameterizations, such as diffusion models. In this paper, we introduce BDPO, a principled behavior-regularized RL framework tailored for diffusion-based policies, thereby combining the expressive power of diffusion policies and the robustness provided by regularization. The key ingredient of our method is to calculate the Kullback-Leibler (KL) regularization analytically as the accumulated discrepancies in reverse-time transition kernels along the diffusion trajectory. By integrating the regularization, we develop an efficient two-time-scale actor-critic RL algorithm that produces the optimal policy while respecting the behavior constraint. Comprehensive evaluations conducted on synthetic 2D tasks and continuous control tasks from the D4RL benchmark validate its effectiveness and superior performance.
@@ 1 Introduction
Despite its huge success in industrial applications such as robotics (Kumar et al., 2021), game AI (Vinyals et al., 2019), and generative model fine-tuning (Ouyang et al., 2022), reinforcement learning (RL) algorithm typically requires millions of online interactions with the environment to achieve meaningful optimization (Haarnoja et al., 2018; Schulman et al., 2017). Given that online interaction can be hazardous and costly in certain scenarios, offline reinforcement learning, which focuses on optimizing policies with static datasets, emerges as a practical and promising avenue (Levine et al., 2020).
@@
Most preliminary works in offline reinforcement learning assume explicit policy distributions (Fujimoto & Gu, 2021; Haarnoja et al., 2018; Christodoulou, 2019); for instance, the agent’s policy is often modeled as a Gaussian distribution with parameterized mean and diagonal covariance, or a deterministic distribution that directly outputs the action. Despite the computational efficiency, such assumptions present significant challenges within the aforementioned behavior-regularized framework (Wang et al., 2023; Hansen-Estruch et al., 2023; Chen et al., 2023). The underlying difficulty originates from the multi-modality of the distributions and manifests in two ways: 1) the dataset may be collected from multiple policy checkpoints, meaning that using unimodal distributions to approximate the behavior policy for regularization can introduce substantial approximation errors (see Figure 1); 2) the optimal policy distribution in behavior-regularized RL framework is actually the Boltzmann distribution of the optimal value function (Haarnoja et al., 2017; Nair et al., 2020), which is inherently multi-modal as well. Such limitation motivates researchers to explore diffusion models, which generate actions through a series of denoising steps, as a viable parameterization of the policy (Wang et al., 2023; Hansen-Estruch et al., 2023; Chen et al., 2024a, b; Fang et al., 2024; Zhang et al., 2024b).
@@
2) Building upon this foundation, we propose a two-time-scale actor-critic method to optimize diffusion policies. Instead of differentiating the policy along the entire diffusion path, BDPO estimates the values at intermediate diffusion steps to amortize the optimization, offering efficient computation, convergence guarantee and state-of-the-art performance;
@@ 2 Related Work
Offline RL. To improve beyond the interaction experience, RL algorithms need to query the estimated values of unseen actions for optimization. In offline scenarios, such distribution shift tends to cause serious overestimation in value functions due to the lack of corrective feedback (Kumar et al., 2020a). To mitigate this, algorithms like CQL (Kumar et al., 2020b), EDAC (An et al., 2021), and PBRL (Bai et al., 2022) focus on penalizing the Q-values of OOD actions to prevent the over-estimation issue. Behavior regularization provides another principled way to mitigate the distribution shift problem, by adding penalties for deviation from the dataset policy during the stage of policy evaluation, policy improvement (Fujimoto et al., 2019; Fujimoto & Gu, 2021; Ran et al., 2023), or sometimes both (Wu et al., 2019; Tarasov et al., 2023). Another line of research, also based on the behavior-regularized RL framework, approaches offline RL by performing in-sample value iteration (Kostrikov et al., 2022; Garg et al., 2023; Xu et al., 2023), thus eliminating OOD queries from the policy and directly approximating the optimal value function. Lastly, model-based offline RL methods (Yu et al., 2020; Sun et al., 2023; Luo et al., 2024) introduce learned dynamics models that generate synthetic experiences to alleviate data limitations, providing extra generalization compared to model-free algorithms.
@@
Diffusion Policies in Offline RL. There has been a notable trend towards the applications of expressive generative models for policy parameterization (Janner et al., 2022), dynamics modeling (Micheli et al., 2023; Alonso et al., 2024), and trajectory planning (Chen et al., 2021; Ajay et al., 2023; Gao et al., 2024). In offline RL, several works employ diffusion models to approximate the behavior policy used for dataset collection. To further improve the policy, they utilize in-sample value iteration to derive the Q-values, and subsequently select action candidates from the behavior diffusion model (Hansen-Estruch et al., 2023; Chen et al., 2023) or use the gradient of Q-value functions to guide the generation (Lu et al., 2023; Mao et al., 2024). However, the performances of these methods are limited, as the actions come from behavior diffusion. Alternatively, SRPO (Chen et al., 2024a) and DTQL (Chen et al., 2024b) maintain a simple one-step policy for optimization while harnessing the diffusion model or diffusion loss to implement behavior regularization. This yields improved computational efficiency and performance in practice; however, the single-step policy still restricts expressiveness and fails to encompass all of the modes of the optimal policy. A wide range of works therefore explore using diffusion models as the actor. Among them, DAC (Fang et al., 2024) formulates the optimization as a noise-regression problem and proposes to align the output from the policy with the gradient of the Q-value functions. Diffusion-QL (Wang et al., 2023) optimizes the actor by back-propagating the gradient of Q-values throughout the entire diffusion path. This results in a significant memory footprint and computational overhead, and EDP (Kang et al., 2023) proposes to use action approximations to alleviate the cost. In contrast, BDPO maintains value functions for intermediate diffusion steps, thereby amortizing the optimization cost while also keeping the optimization precise.
@@ 3 Preliminaries
Behavior-Regularized Offline Reinforcement Learning. We formalize the task as a Markov Decision Process (MDP) $\langle\mathcal{S},\mathcal{A},T,R,\gamma\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $T(s^{\prime}|s,a)$ denotes the transition function, $R(s,a)$ is a bounded reward function, and $\gamma$ is the discount factor. In offline RL, an agent is expected to learn a policy $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ to maximize the expected discounted return $\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})]$ with an offline dataset $\mathcal{D}=\{(s_{t},a_{t},s_{t+1},r_{t})\}$, where $s_{t},s_{t+1}\in\mathcal{S}$, $a_{t}\in\mathcal{A}$, and $r_{t}=R(s_{t},a_{t})\in\mathbb{R}$. We consider the behavior-regularized RL objective, which augments the original RL objective by regularizing the policy towards some behavior policy $\nu$:
@@
where $\eta>0$ controls the regularization strength. In offline RL, Eq. (1) is widely employed by setting $\nu$ as the policy $\pi_{\mathcal{D}}$ that collects the dataset to prevent the exploitation of the out-of-dataset actions. Besides, when setting $\nu$ as the uniform distribution, Eq. (1) equates to the maximum-entropy RL in online scenarios up to some constant.
@@
where $s^{\prime}\sim T(\cdot|s,a)$ and $a^{\prime}\sim\pi(\cdot|s^{\prime})$. For policy improvement, we can update the policy as follows:
$\max_{\pi}\ \mathbb{E}_{a\sim\pi(a|s)}\left[Q^{\pi}(s,a)\right]-\eta\operatorname{KL}\left[\pi(\cdot|s)\|\nu(\cdot|s)\right].$ (4)
@@
Diffusion Models. Diffusion models (Ho et al., 2020; Song et al., 2021) consist of a forward Markov process which progressively perturbs the data $x^{0}\sim q_{0}$ to data that follows the standard Gaussian distribution $x^{N}\sim q_{N}(x)$, and a reverse Markov process that gradually recovers the original $x^{0}$ from the noisy sample $x^{N}$. The transition of the forward process $q_{n+1|n}$ usually follows Gaussian distributions:
@@
where $\{\beta_{n}\}_{n=1}^{N}$ is specified according to the noise schedule. Due to the closure property of Gaussian distributions, the marginal distribution of $x^{n}$ given $x^{0}$ can be specified as:
@@
where $I$ is the identity matrix, and $\sigma_{n}=\sqrt{\frac{1-\bar{\alpha}_{n-1}}{1-\bar{\alpha}_{n}}\beta_{n}}$. The learning objective is matching $p^{\theta}_{n-1|n}(x^{n-1}|x^{n})$ with the posterior $q_{n-1|n,0}(x^{n-1}|x^{n},x^{0})$:
$\mathcal{L}_{\rm diff}(\theta)=$ (9)
$\mathbb{E}_{n,x^{0},x^{n}}\left[\operatorname{KL}\left[q_{n-1|n,0}(x^{n-1}|x^{n},x^{0})\|p^{\theta}_{n-1|n}(x^{n-1}|x^{n})\right]\right]$
where $x^{0}\sim q_{0},x^{n}\sim q_{n|0}(\cdot|x^{0})$. Simplified training objectives can be derived by reparameterizing $\mu^{\theta}$ with noise prediction or score networks (Ho et al., 2020; Song et al., 2021). After training, the generation process begins by sampling $\hat{x}_{N}\sim\mathcal{N}(0,I)$, followed by iteratively applying $p^{\theta}_{n-1|n}$ to generate the final samples $\hat{x}_{0}$, which approximately follow the target distribution $q_{0}$.
@@ 4.1 Pathwise KL Regularization
In behavior-regularized RL, previous methods typically regularize action distribution, i.e. the marginalized distribution $p^{\pi,s}_{0}$. Instead, we shift our attention to the KL divergence with respect to the diffusion path $a^{0:N}$, which can be further decomposed thanks to the Markov property:
$\operatorname{KL}\left[p^{\pi,s}_{0:N}\|p^{\nu,s}_{0:N}\right]$ (10)
@@
$=\mathbb{E}\left[\sum_{n=1}^{N}\operatorname{KL}\left[p^{\pi,s,a^{n}}_{n-1|n}\|p^{\nu,s,a^{n}}_{n-1|n}\right]\right].$
Here, the expectation is taken w.r.t. $a^{0:N}\sim p^{\pi,s}_{0:N}$, $p^{\nu}$ is the diffusion policy trained via Eq. (9) on the offline dataset $\mathcal{D}$ to approximate the data-collection policy, and the last equation holds due to $p_{N}^{\pi,s}=p_{N}^{\nu,s}=\mathcal{N}(0,I)$. In the following, we abbreviate $\operatorname{KL}\left[p^{\pi,s,a^{n}}_{n-1|n}\|p^{\nu,s,a^{n}}_{n-1|n}\right]$ with $\ell^{\pi,s}_{n}(a^{n})$.
@@
(Pathwise KL-Regularized RL) Let $p^{\nu}$ be the reference diffusion process. The pathwise KL-regularized RL problem seeks to maximize
$\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(R(s_{t},a_{t}^{0})-\eta\sum_{n=1}^{N}\ell_{n}^{\pi,s_{t}}(a_{t}^{n})\right)\right].$ (11)
Although our problem calculates the accumulated discrepancies along the diffusion path as penalties, it actually preserves the same optimal solution to Eq. (1) which only regularizes the action distribution at diffusion step $n=0$. The equivalency is captured by the following theorem.
@@
Let $p^{\nu}$ be the reference behavior diffusion process. The optimal diffusion policy $p^{*}$ of the pathwise KL-regularized RL in Eq. (11) is also the optimal policy $\pi^{*}$ of the KL regularized objective in Eq. (1) in the sense that $\pi^{*}(a|s)=\int p^{*,s}_{0:N}(a^{0:N})\delta(a-a^{0})\mathrm{d}a^{0:N}\forall s\in\mathcal{S}$, where $\delta$ is the Dirac delta function.
Proof.
Proof can be found in Appendix C.2. ∎
@@
Figure 3: Semantic illustration of the TD backup for the Q-value function $Q^{\pi}$ (blue) and diffusion value function $V^{\pi,s}_{n}$ (orange). The update of $Q^{\pi}$ (Eq. (12)) requires reward, penalties along the diffusion trajectory, and the Q-values at next state. The update of $V^{\pi,s}_{n}$ (Eq. (14)) involves the single-step penalty and the diffusion value at the next diffusion step $n-1$.

### v2

@@ Abstract
Behavior regularization, which constrains the policy to stay close to some behavior policy, is widely used in offline reinforcement learning (RL) to manage the risk of hazardous exploitation of unseen actions. Nevertheless, existing literature on behavior-regularized RL primarily focuses on explicit policy parameterizations, such as Gaussian policies. Consequently, it remains unclear how to extend this framework to more advanced policy parameterizations, such as diffusion models. In this paper, we introduce BDPO, a principled behavior-regularized RL framework tailored for diffusion-based policies, thereby combining the expressive power of diffusion policies and the robustness provided by regularization. The key ingredient of our method is to calculate the Kullback-Leibler (KL) regularization analytically as the accumulated discrepancies in reverse-time transition kernels along the diffusion trajectory. By integrating the regularization, we develop an efficient two-time-scale actor-critic RL algorithm that produces the optimal policy while respecting the behavior constraint. Comprehensive evaluations conducted on synthetic 2D tasks and continuous control tasks from the D4RL benchmark validate its effectiveness and superior performance. The code and experiment results of BDPO are available on the project webpage.
@@ 1 Introduction
Despite its huge success in industrial applications such as robotics (Kumar et al., 2021), game AI (Vinyals et al., 2019), and generative model fine-tuning (Ouyang et al., 2022), reinforcement learning (RL) algorithms typically require millions of online interactions with the environment to achieve meaningful optimization (Haarnoja et al., 2018; Schulman et al., 2017). Given that online interaction can be hazardous and costly in certain scenarios, offline RL, which focuses on optimizing policies with static datasets, emerges as a practical and promising avenue (Levine et al., 2020).
@@
Most preliminary works in offline RL assume explicit policy distributions (Fujimoto & Gu, 2021; Haarnoja et al., 2018; Christodoulou, 2019); for instance, the agent’s policy is often modeled as a Gaussian distribution with parameterized mean and diagonal covariance, or a deterministic distribution that directly outputs the action. Despite the computational efficiency, such assumptions present significant challenges within the aforementioned behavior-regularized framework (Wang et al., 2023; Hansen-Estruch et al., 2023; Chen et al., 2023). The underlying difficulty originates from the multi-modality of the distributions and manifests in two ways: 1) the dataset may be collected from multiple policy checkpoints, meaning that using unimodal distributions to approximate the behavior policy for regularization can introduce substantial approximation errors (see Figure 1); 2) the optimal policy distribution in behavior-regularized RL framework is actually the Boltzmann distribution of the optimal value function (Haarnoja et al., 2017; Nair et al., 2020), which is inherently multi-modal as well. Such limitation motivates researchers to explore diffusion models, which generate actions through a series of denoising steps, as a viable parameterization of the policy (Wang et al., 2023; Hansen-Estruch et al., 2023; Chen et al., 2024a, b; Fang et al., 2024; Zhang et al., 2024b).
@@
2) Building upon this foundation, we propose a two-time-scale actor-critic method to optimize diffusion policies. Instead of differentiating the policy along the entire diffusion path, BDPO estimates the values at intermediate diffusion steps to amortize the optimization, offering efficient computation, convergence guarantee, and state-of-the-art performance;
@@ 2 Related Work
Offline RL. To improve beyond the interaction experience, RL algorithms need to query the estimated values of unseen actions for optimization. In offline scenarios, such distribution shift tends to cause serious overestimation in value functions due to the lack of corrective feedback (Kumar et al., 2020a). To mitigate this, algorithms like CQL (Kumar et al., 2020b), EDAC (An et al., 2021), and PBRL (Bai et al., 2022) focus on penalizing the Q-values of OOD actions to prevent the over-estimation issue. Behavior regularization provides another principled way to mitigate the distribution shift problem, by adding penalties for deviation from the dataset policy during the stage of policy evaluation, policy improvement (Fujimoto et al., 2019; Fujimoto & Gu, 2021; Ran et al., 2023), or sometimes both (Wu et al., 2019; Tarasov et al., 2023). Another line of research, also based on the behavior-regularized RL framework, approaches offline RL by performing in-sample value iteration (Kostrikov et al., 2022; Garg et al., 2023; Xu et al., 2023), thus eliminating OOD queries from the policy and directly approximating the optimal value function. Lastly, model-based offline RL methods (Yu et al., 2020; Jia et al., 2024; Sun et al., 2023; Luo et al., 2024) introduce learned dynamics models that generate synthetic experiences to alleviate data limitations, providing extra generalization compared to model-free algorithms.
@@
Diffusion Policies in Offline RL. There has been a notable trend towards the applications of expressive generative models for policy parameterization (Janner et al., 2022), dynamics modeling (Micheli et al., 2023; Alonso et al., 2024), trajectory planning (Chen et al., 2021; Ajay et al., 2023; Gao et al., 2024), and representation learning (Shribak et al., 2024). In offline RL, several works employ diffusion models to approximate the behavior policy used for dataset collection. To further improve the policy, they utilize in-sample value iteration to derive the Q-values, and subsequently select action candidates from the behavior diffusion model (Hansen-Estruch et al., 2023; Chen et al., 2023) or use the gradient of Q-value functions to guide the generation (Lu et al., 2023; Mao et al., 2024). However, the performance of these methods is limited, as the actions come from behavior diffusion. Alternatively, SRPO (Chen et al., 2024a) and DTQL (Chen et al., 2024b) maintain a simple one-step policy for optimization while harnessing the diffusion model or diffusion loss to implement behavior regularization. This yields improved computational efficiency and performance in practice; however, the single-step policy still restricts expressiveness and fails to encompass all of the modes of the optimal policy. A wide range of works therefore explore using diffusion models as the actor. Among them, DAC (Fang et al., 2024) formulates the optimization as a noise-regression problem and proposes to align the output from the policy with the gradient of the Q-value functions. Diffusion-QL (Wang et al., 2023) optimizes the actor by back-propagating the gradient of Q-values throughout the entire diffusion path. This results in a significant memory footprint and computational overhead, and EDP (Kang et al., 2023) proposes to use action approximations to alleviate the cost. In contrast, BDPO maintains value functions for intermediate diffusion steps, thereby amortizing the optimization cost while also keeping the optimization precise.
@@ 3 Preliminaries
Behavior-Regularized Offline RL. We formalize the task as a Markov Decision Process (MDP) $\langle\mathcal{S},\mathcal{A},T,R,\gamma\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $T(s^{\prime}|s,a)$ denotes the transition function, $R(s,a)$ is a bounded reward function, and $\gamma$ is the discount factor. In offline RL, an agent is expected to learn a policy $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ to maximize the expected discounted return $\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})]$ with an offline dataset $\mathcal{D}=\{(s_{t},a_{t},s_{t+1},r_{t})\}$, where $s_{t},s_{t+1}\in\mathcal{S}$, $a_{t}\in\mathcal{A}$, and $r_{t}=R(s_{t},a_{t})\in\mathbb{R}$. We consider the behavior-regularized RL objective, which augments the original RL objective by regularizing the policy towards some behavior policy $\nu$:
@@
where $D_{\rm KL}$ is the Kullback-Leibler (KL) divergence and $\eta>0$ controls the regularization strength. In offline RL, Eq. (1) is widely employed by setting $\nu$ as the policy $\pi_{\mathcal{D}}$ that collects the dataset to prevent the exploitation of the out-of-dataset actions. Besides, when setting $\nu$ as the uniform distribution, Eq. (1) equates to the maximum-entropy RL in online scenarios up to some constant.
@@
where $s^{\prime}\sim T(\cdot|s,a)$ and $a^{\prime}\sim\pi(\cdot|s^{\prime})$. For policy improvement, we can update the policy using the objective:
$\max_{\pi}\ \mathbb{E}_{a\sim\pi(\cdot|s)}\left[Q^{\pi}(s,a)\right]-\eta D_{\rm KL}\left[\pi(\cdot|s)\|\nu(\cdot|s)\right].$ (4)
@@
Diffusion Models. Diffusion models (Ho et al., 2020; Song et al., 2021) consist of a forward Markov process which progressively perturbs the data $x^{0}\sim q_{0}$ to data that approximately follows the standard Gaussian distribution $x^{N}\sim q_{N}$, and a reverse Markov process that gradually recovers the original $x^{0}$ from the noisy sample $x^{N}$. The transition of the forward process $q_{n+1|n}$ usually follows Gaussian distributions:
@@
where $\{\beta_{n}\}_{n=1}^{N}$ is specified according to the noise schedule and $I$ denotes the Identity matrix. Due to the closure property of Gaussian distributions, the marginal distribution of $x^{n}$ given $x^{0}$ can be specified as:
@@
where $\sigma_{n}=\sqrt{\frac{1-\bar{\alpha}_{n-1}}{1-\bar{\alpha}_{n}}\beta_{n}}\approx\sqrt{\beta_{n}}$. The learning objective is matching $p^{\theta}_{n-1|n}(x^{n-1}|x^{n})$ with the posterior $q_{n-1|n,0}(x^{n-1}|x^{n},x^{0})$:
[…]
where $x^{0}\sim q_{0},x^{n}\sim q_{n|0}(\cdot|x^{0})$. Upon the completion of training, it can be shown that the objective Eq. (9) yields
$p^{\theta}_{n-1|n}(x^{n-1}|x^{n})\approx q_{n-1|n}(x^{n-1}|x^{n}).$ (10)
Simplified training objectives can be derived by reparameterizing $\mu^{\theta}$ with noise prediction or score networks (Ho et al., 2020; Song et al., 2021). After training, the generation process begins by sampling $\hat{x}^{N}\sim\mathcal{N}(0,I)$, followed by iteratively applying $p^{\theta}_{n-1|n}$ to generate the final samples $\hat{x}_{0}$, which approximately follow the target distribution $q_{0}$.
@@ 4.1 Pathwise KL Regularization
In behavior-regularized RL, previous methods typically regularize action distribution, i.e., the marginalized distribution $p^{\pi,s}_{0}$. Instead, we shift our attention to the KL divergence with respect to the diffusion path $a^{0:N}$, which can be further decomposed thanks to the Markov property:
$D_{\rm KL}\left[p^{\pi,s}_{0:N}\|p^{\nu,s}_{0:N}\right]$ (11)
@@
$=\mathbb{E}\left[\sum_{n=1}^{N}D_{\rm KL}\left[p^{\pi,s,a^{n}}_{n-1|n}\|p^{\nu,s,a^{n}}_{n-1|n}\right]\right].$
Here, the expectation is taken w.r.t. $a^{0:N}\sim p^{\pi,s}_{0:N}$, $p^{\nu}$ is the behavior diffusion policy trained via Eq. (9) on the offline dataset $\mathcal{D}$ to approximate the data-collection policy, and the last equation holds due to $p_{N}^{\pi,s}=p_{N}^{\nu,s}=\mathcal{N}(0,I)$. In the following, we abbreviate $D_{\rm KL}\left[p^{\pi,s,a^{n}}_{n-1|n}\|p^{\nu,s,a^{n}}_{n-1|n}\right]$ with $\ell^{\pi,s}_{n}(a^{n})$.
@@
(Pathwise KL-Regularized RL) Let $p^{\nu}$ be the behavior diffusion process. The pathwise KL-regularized RL problem seeks to maximize the following objective
$\max_{p^{\pi}}\ \mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(R(s_{t},a_{t}^{0})-\eta\sum_{n=1}^{N}\ell_{n}^{\pi,s_{t}}(a_{t}^{n})\right)\right].$ (12)
where the expectation is taken w.r.t. trajectories following $p^{\pi}$ and the dynamics $T$.
Although our problem calculates the accumulated discrepancies along the diffusion path as penalties, it actually preserves the same optimal solution to Eq. (1), which only regularizes the action distribution at diffusion step $n=0$. The following theorem captures the equivalence.
@@
(Proof in Appendix C.2) Let $p^{\nu}$ be the behavior diffusion process. The optimal diffusion policy $p^{*}$ of the pathwise KL-regularized RL problem in Eq. (12) is also the optimal policy $\pi^{*}$ of the KL regularized objective in Eq. (1), in the sense that $\pi^{*}(a|s)=\int p^{*,s}_{0:N}(a^{0:N})\delta(a-a^{0})\mathrm{d}a^{0:N}\ \forall s\in\mathcal{S}$, where $\delta$ is the Dirac delta function.
@@
Figure 3: Semantic illustration of the TD backup for the Q-value function $Q^{\pi}$ (blue) and diffusion value function $V^{\pi,s}_{n}$ (orange). The update of $Q^{\pi}$ (Eq. (13)) requires reward, penalties along the diffusion trajectory, and the Q-values at the next state. The update of $V^{\pi,s}_{n}$ (Eq. (15)) involves the single-step penalty and the diffusion value at the next diffusion step $n-1$.